# Convergence of a Finite Volume Scheme for Compactly Heterogeneous Scalar Conservation Laws

### v2

@@ Abstract
11footnotetext: abraham.sylla@unibs.it
Department of Information Engineering, University of Brescia (Italy)
ORCID number: $0$-0003-1784-4878
@@
Keywords: Scalar Conservation Law, Compactly Space Dependent Flux, Finite Volume Scheme, Discontinuous Flux, Well-Balanced Scheme.
@@ 1 Introduction
Equations of this type often occur, for instance, in the modeling of physical phenomena related to traffic flow [40, 31, 20], porous media [19, 21] or sedimentation problems [13, 14].
It is known that solutions of (1) are discontinuous, regardless of the regularity of the data. They are to be understood in the entropy sense of [27]. The following quantity often recurs below,
@@
Let $H\in\mathbf{C}^{1}(\mathbb{R}^{2},\mathbb{R})$ and $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. We say that $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ is an entropy solution to (1) if for all $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and $k\in\mathbb{R}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}|u-k|\partial_{t}\varphi+\Phi(x,u,k)\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}-\int_{0}^{+\infty}\int_{\mathbb{R}}\mathop{\rm sgn}(u-k)\partial_{x}H(x,k)\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$+\int_{\mathbb{R}}|u_{o}(x)-k|\varphi(0,x)\mathinner{\mathrm{d}{x}}\geq 0.$ (1.2)
In appearance, Definition 1.1 is weaker than the classical [27, Definition 1] since it does not require the existence of a strong trace at the initial time. It is in particular more manageable to limit process. Nevertheless, it can be shown that if $H\in\mathbf{C}^{3}(\mathbb{R}^{2},\mathbb{R})$, then Definition 1.1 ensures that each entropy solution admits a representative belonging to $\mathbf{C}(\mathbb{R}^{+},\mathbf{L}_{\mathbf{loc}}^{1}(\mathbb{R},\mathbb{R}))$, see [11, Theorem 2.6].
In [27], the author proved existence, uniqueness and stability with the respect to the initial datum for (1) in the framework of entropy solutions under, among others, the growth assumptions
$\partial_{u}H\in\mathbf{L}^{\infty}(\mathbb{R}^{2},\mathbb{R})\quad\text{and}\quad\sup_{(x,u)\in\mathbb{R}^{2}}\left(-\partial_{xu}^{2}H(x,u)\right)<+\infty,$ (1.3)
highlighting that the space dependency in (1) was treated as a source term. In his paper, Kruzhkov proved existence using the vanishing viscosity method, that is, through a parabolic regularization of (1). The following example motivates the necessity to relax (1.3) in the 1D scalar case.
@@
Fix positive constants $X,V_{1},V_{2},R_{1},R_{2}$, and let $\mathcal{V},\rho\in\mathbf{C}^{3}(\mathbb{R},\mathopen{]}0,+\infty\mathclose{[})$ be such that $\mathcal{V}(x)=V_{1}$, resp. $\rho(x)=R_{1}$, for $x<-X$ and $\mathcal{V}(x)=V_{2}$, resp. $\rho(x)=R_{2}$, for $x>X$. Define
@@
Then with this flow, (1) is the so-called “LWR” (Lightill-Whitham, Richards) model [28, 34] for a flow of vehicles described by their density $u$ along a rectilinear road with maximal speed, resp. density, smoothly varying from $V_{1}$ to $V_{2}$, resp. from $R_{1}$ to $R_{2}$. Clearly, $H$ does not satisfy (1.3).
Remark 1.1.
A priori, Kruzhkov results do not apply in the case of Example 1.1. For completeness, let us however mention that a truncation argument could be used to extend them if the initial datum takes values between the stationary solutions $u(t,x)=0$ and $u(t,x)=\rho(x)$.
Recently, the authors of [11] developed an alternative framework to tackle (1), one not requiring (1.3) and inspired by Example 1.1:
@@
$\mathbf{Compact\leavevmode\nobreak\ Heterogeneity:}$ $\quad\exists X>0,\;\forall(x,u)\in\mathbb{R}^{2},\quad|x|\geq X\implies\partial_{u}H(x,u)=0.$ (CH)
$\mathbf{Strong\leavevmode\nobreak\ Convexity:}$ $\quad\forall x\in\mathbb{R},\;u\mapsto\partial_{u}H(x,u)\;\text{is an increasing}$ (CVX)
@@
Like we previously mentioned, Assumptions (C3)–(CVX) comprise flows that do not fit in the classical Kruzhkov framework and that are relevant from the modeling point of view, for instance, the flow of Example 1.1 (in the concave case). On the other hand, Kruzhkov results apply to general balance laws in several space dimensions. Let us recall the following.
@@
[11, Theorem 2.6 and Theorem 2.18] Assume that $H$ satisfies (C3)–(CVX). Then for all $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$, the Cauchy problem (1) admits a unique entropy solution $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
@@
$\int_{|x|\leq R}|u(t,x)-v(t,x)|\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}|u_{o}(x)-v_{o}(x)|\mathinner{\mathrm{d}{x}}$
$\int_{|x|\leq R}(u(t,x)-v(t,x))^{+}\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}(u_{o}(x)-v_{o}(x))^{+}\mathinner{\mathrm{d}{x}}.$
The proof of existence in [11], like in [27], relies on the vanishing viscosity method, but there, the authors exploit the correspondence with the Hamilton-Jacobi equation (and its parabolic approximation), going back and forth between the two frameworks, and gathering information for both equations at each step.
@@
To be precise, let us mention that in [11], the convexity assumption (CVX) is relaxed to a uniform coercivity assumption coupled to a genuine nonlinearity assumption:
$\mathbf{Uniform\leavevmode\nobreak\ Coercivity:}$ $\quad\forall h\in\mathbb{R},\;\exists\;\mathcal{U}_{h}\in\mathbb{R},\;\forall(x,u)\in\mathbb{R}^{2},$
$\quad|H(x,u)|\leq h\implies|u|\leq\mathcal{U}_{h}.$ (UC)
[…]
@@
[…]
$\forall x,u\in\mathbb{R},\;H(x,u)\geq\phi(|u|)\quad\text{and}\quad\frac{\phi(r)}{r}{\ \underset{r\to+\infty}{\longrightarrow}\ }+\infty.$
The convexity was used in [12] to characterize, for (1), the attainable set and the set of initial data evolving at a prescribed time into a prescribed profile.
An alternative to the vanishing viscosity method to construct solutions to (1) is to build and prove the convergence of finite volume schemes. The most recent results include the works of [18, 17, 9] (in the case $H(x,u)=\theta(x)f(u)$), [7] (convergence and error analysis), [8] (with a source term) or [39] (on a bounded domain). In all these works, the flux function enters the framework of Kruzhkov by satisfying either (1.3), or the stronger requirement $\partial_{x}H\equiv 0$.
The aim of this paper is to build a finite volume scheme for (1) and prove its convergence to the entropy solution under the assumptions (C3)–(CVX). As a byproduct, we provide an alternate existence result for the Cauchy problem (1), one that does not rely on the vanishing viscosity method. However, let us stress that the main result of the paper is a convergence result and not an existence result. To the author’s knowledge, no convergence result is available for flux functions that do not verify (1.3).
In the convergence analysis of any standard finite volume scheme for (1), a difficulty to tackle is the obtaining of a priori $\mathbf{L}^{\infty}$ bounds for the approximating sequence. Indeed, because of the space dependency, constants are not steady states anymore, therefore the maximum principle breaks down. In [18, 17, 9, 7, 8, 39], this issue was overcome either with the requirement $\partial_{x}H\equiv 0$, ensuring that constants are steady states, or with Assumption (1.3), which provides a time dependent $\mathbf{L}^{\infty}$ bounds of the numerical solution. We choose a new approach, relying on the theory of discontinuous flux, see [1, 2, 3, 4, 13, 19, 24, 26, 30, 37] and the references therein. For the construction of the scheme, the space dependency of $H$ is discretized, and the idea is to treat each edge of the mesh as an interface of a discontinuous flux problem. Outside the compact of space dependency, our scheme reduces to a standard Godunov method. The contribution of the discontinuous flux theory is that we can build by hand discrete steady states of the scheme, making it, in a way well-balanced.
@@
2 Description of the scheme and main results
@@
$\forall j\in\mathbb{Z},\quad u_{j}^{o}\coloneq\frac{1}{\Delta x}\int_{I_{j}}u_{o}(x)\mathinner{\mathrm{d}{x}},\quad w_{\Delta}(x)\coloneq\sum_{j\in\mathbb{Z}}x_{j}\mathbf{1}_{I_{j}}(x).$
Fix $n\in\mathbb{N}$. Let us explain how, given $(u_{j}^{n})_{j\in\mathbb{Z}}$, we determine $(u_{j}^{n+1})_{j\in\mathbb{Z}}$. Set $\rho^{n}\coloneq\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{I_{j}}$. Let us call $\mathcal{U}^{n}=\mathcal{U}^{n}(t,x)$ the unique entropy solution to following Cauchy problem, set on $\mathopen{]}t^{n},t^{n+1}\mathopen{[}\times\mathbb{R}$:
$\partial_{t}\mathcal{U}(t,x)+\partial_{x}\bigl{(}H(w_{\Delta}(x),\mathcal{U}(t,x))\bigr{)}$ $=0$
$\mathcal{U}(t^{n},x)$ $=\rho^{n}(x).$ (2.1)
We recall in Section 3.1 the meaning of solutions to conservation laws with discontinuous flux we use and also establish that (2) admits a unique solution. Notice that since the space dependency is localized in a compact subset of $\mathbb{R}$, outside this compact, (2) reduces to a standard homogeneous conservation law.
Define $u_{j}^{n+1}$ as
$\forall j\in\mathbb{Z},\quad u_{j}^{n+1}\coloneq\frac{1}{\Delta x}\int_{I_{j}}\mathcal{U}^{n}(t^{n+1}-,x)\mathinner{\mathrm{d}{x}}.$ (2.2)
@@
Assume that $H$ satisfies (C3)–(CVX) and fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. There exist two constants $\underline{u}$, $\overline{u}$ depending only on $H$ and $u_{o}$ such that under the CFL condition
$2\lambda L\leq 1,\quad\lambda\coloneq\frac{\Delta t}{\Delta x},\;L\coloneq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}|\partial_{u}H(x,p)|,$ (2.3)
@@
[…]
@@
The proof of (M2) is then somewhat standard, and the details are in Section 3.3. For the convergence analysis, since global $\mathbf{BV}$ bounds are not expected in the context of discontinuous flux, we rely on the compensated compactness method [29, 35], taking advantage of the genuine nonlinearity of $H$ under Assumption (CVX).

### v3

@@ Abstract
11footnotetext: abraham.sylla@u-picardie.fr
LAMFA CNRS UMR 7352, Université de Picardie Jules Verne
33 rue Saint-Leu, 80039 Amiens (France)
ORCID number: 0000-0003-1784-4878
@@
Keywords: Scalar Conservation Law, Compactly Space Dependent Flux, Discontinuous Flux, Finite Volume Scheme, Well-Balanced Scheme.
@@ 1 Introduction
Equations of this type often occur, for instance, in the modeling of physical phenomena related to traffic flow [Whitham2014, Mochon1987, HR1995], porous media [GR1992, Jaffre1995] or sedimentation problems [Diehl1995, Diehl1996].
It is known that solutions of (1) are discontinuous, regardless of the regularity of the data. They are to be understood in the entropic sense of [Kruzhkov1970]. The following quantity often recurs below,
@@
Let $H\in\mathbf{C}^{1}(\mathbb{R}^{2},\mathbb{R})$ and $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. We say that $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$ is an entropy solution to (1) if for all test functions $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and $k\in\mathbb{R}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}{\left|u-k\right|}\partial_{t}\varphi+\Phi(x,u,k)\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}-\int_{0}^{+\infty}\int_{\mathbb{R}}\mathop{\rm sgn}(u-k)\partial_{x}H(x,k)\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}$
$+\int_{\mathbb{R}}{\left|u_{o}(x)-k\right|}\varphi(0,x)\mathinner{\mathrm{d}{x}}\geq 0.$ (1.2)
In appearance, Definition 1.1 is weaker than the classical [Kruzhkov1970, Definition 1] since it does not require the existence of a strong trace at the initial time. It is in particular more manageable to limit process. Nevertheless, it can be shown that if $H\in\mathbf{C}^{3}(\mathbb{R}^{2},\mathbb{R})$, then Definition 1.1 ensures that each entropy solution admits a representative belonging to $\mathbf{C}(\mathbb{R}^{+},\mathbf{L}_{\mathbf{loc}}^{1}(\mathbb{R},\mathbb{R}))$, see [CPS2023, Theorem 2.6].
In [Kruzhkov1970], the author proved existence, uniqueness and stability with respect to the initial datum for (1) in the framework of entropy solutions under, among others, the growth assumptions
$\partial_{u}H\in\mathbf{L}^{\infty}(\mathbb{R}^{2},\mathbb{R})\quad\text{and}\quad\sup_{(x,u)\in\mathbb{R}^{2}}\left(-\partial_{xu}^{2}H(x,u)\right)<+\infty.$ (1.3)
These assumptions highlight the fact that the space dependency of $H$ in (1) was treated as a source term. In his paper, Kruzhkov proved existence using the vanishing viscosity method, that is, through a parabolic regularization of (1). The following example motivates the necessity to relax (1.3) in the 1D scalar case.
@@
Fix positive constants $X,V_{1},V_{2},R_{1},R_{2}$, and let $\mathcal{V},\rho\in\mathbf{C}^{3}(\mathbb{R},\mathopen{]}0,+\infty\mathclose{[})$ be such that $\mathcal{V}(x)=V_{1}$, resp. $\rho(x)=R_{1}$, for x<Xx<-X and $\mathcal{V}(x)=V_{2}$, resp. $\rho(x)=R_{2}$, for $x>X$. Define
@@
Then with this flow, (1) is the so-called “LWR” (Lightill-Whitham, Richards) model [LW1955, Richards1956] for a flow of vehicles described by their density $u$ along a rectilinear road with maximal speed, resp. density, smoothly varying from $V_{1}$ to $V_{2}$, resp. from $R_{1}$ to $R_{2}$. Clearly, $H$ does not satisfy (1.3). Let us signal however, that since 0 and $\rho$ are stationary solutions to the equation, this example could treated in Kruzhkov’s framework for nonnegative initial data bounded from above by $\rho$.
Recently, the authors of [CPS2023] developed an alternative framework to tackle (1), one not requiring the growth assumptions (1.3) and inspired by Example 1.1:
@@
$\mathbf{Compact\penalty 10000\ Heterogeneity:}$ $\quad\exists X>0,\;\forall(x,u)\in\mathbb{R}^{2},\quad{\left|x\right|}\geq X\implies\partial_{x}H(x,u)=0.$ (CH)
$\mathbf{Strong\penalty 10000\ Convexity:}$ $\quad\forall x\in\mathbb{R},\;u\mapsto\partial_{u}H(x,u)\;\text{is an increasing}$ (CVX)
@@
Like we previously mentioned, Assumptions (C3)–(CH)–(CVX) comprise flows that do not fit in the classical Kruzhkov framework and that are relevant from the modeling point of view, for instance, the flow of Example 1.1 (in the concave case). Other examples are given by flows of the type
$H(x,u)\coloneq\theta(x)h\bigl(u-\ell(x)\bigr)+g(x),$
for suitable assumptions on the functions $\theta,h,\ell$ and $g$. More precisely, $\theta,\ell,g\in\mathbf{C}^{3}(\mathbb{R},\mathbb{R})$ have derivatives with compact support, $\theta>0$ and $h\in\mathbf{C}^{3}(\mathbb{R},\mathbb{R})$ is such that $h^{\prime}$ is a $\mathbf{C}^{1}$-diffeomorphism from $\mathbb{R}$ onto itself. A priori, those flows cannot be treated with a truncation argument.
On the other hand, Kruzhkov results apply to general balance laws in several space dimensions. Let us recall the following.
@@
[CPS2023, Theorem 2.6 and Theorem 2.18] Assume that $H$ satisfies (C3)–(CH)–(CVX). Then for all $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$, the Cauchy problem (1) admits a unique entropy solution $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$.
@@
$\int_{{\left|x\right|}\leq R}{\left|u(t,x)-v(t,x)\right|}\mathinner{\mathrm{d}{x}}$ $\leq\int_{{\left|x\right|}\leq R+Lt}{\left|u_{o}(x)-v_{o}(x)\right|}\mathinner{\mathrm{d}{x}}$
The proof of existence in [CPS2023], like in [Kruzhkov1970], relies on the vanishing viscosity method, but there, the authors exploit the correspondence with the Hamilton-Jacobi equation (and its parabolic approximation), going back and forth between the two frameworks, and gathering information for both equations at each step.
@@
To be precise, let us mention that in [CPS2023], the convexity assumption (CVX) is relaxed to a uniform coercivity assumption coupled to a genuine nonlinearity assumption:
$\mathbf{Uniform\penalty 10000\ Coercivity:}$ $\quad\forall h\in\mathbb{R},\;\exists\;\mathcal{U}_{h}\in\mathbb{R},\;\forall(x,u)\in\mathbb{R}^{2},$
$\quad{\left|H(x,u)\right|}\leq h\implies{\left|u\right|}\leq\mathcal{U}_{h}.$ (UC)
$\mathbf{Weak\penalty 10000\ Genuine\penalty 10000\ NonLinearity:}$ $\quad\text{for a.e.}\;x\in\mathbb{R},\;\text{the set}\;\{p\in\mathbb{R}\;:\;\partial_{uu}^{2}H(x,p)=0\}$
@@
The strong convexity implies (1.1), while (1.1) follows from the fact that $H$ admits a Nagumo function (see [CannarsaSinestrariSBook, Theorem 6.1.2]), meaning that exists a function $\phi\in\mathbf{C}(\mathbb{R}^{+},\mathbb{R})$ such that:
$\forall x,u\in\mathbb{R},\;H(x,u)\geq\phi({\left|u\right|})\quad\text{and}\quad\frac{\phi(r)}{r}{\ \underset{r\to+\infty}{\longrightarrow}\ }+\infty.$
For instance,
$\forall r\in\mathbb{R}^{+},\quad\phi(r)\coloneq\inf_{x\in\mathbb{R}}\bigl(\min\{H(x,r),H(x,-r)\}\bigr)$
is such a function.
The convexity was used in [CPSInverseDesign] to characterize, for (1), the attainable set and the set of initial data evolving at a prescribed time into a prescribed profile.
To construct solutions of (1), an alternative to the vanishing viscosity is to build and prove the convergence of finite volume schemes. The most recent results include the works of [EGH1995, EGGH1998, Champier1999] (in the case $H(x,u)=\theta(x)f(u)$), [CH1999] (convergence and error analysis), [CHC2001] (with a source term) or [Vovelle2002] (on a bounded domain). In all these works, the flux function enters the framework of Kruzhkov by satisfying either (1.3), or the stronger requirement $\partial_{x}H\equiv 0$.
The aim of this paper is to build a finite volume scheme for (1) and prove its convergence to the entropy solution under the assumptions (C3)–(CH)–(CVX). As a byproduct, we provide an alternate existence result for the Cauchy problem (1), one that does not rely on the vanishing viscosity method. However, let us stress that the main result of the paper is a convergence result and not an existence result. To the author’s knowledge, no convergence result is available for flux functions that do not verify (1.3).
In the convergence analysis of any standard finite volume scheme for (1), a difficulty to tackle is the obtaining of a priori $\mathbf{L}^{\infty}$ bounds for the approximating sequence. Indeed, because of the space dependency, constants are not steady states anymore, therefore the maximum principle breaks down. In [EGH1995, EGGH1998, Champier1999, CH1999, CHC2001, Vovelle2002], this issue is overcome either with the requirement $\partial_{x}H\equiv 0$, ensuring that constants are steady states, or with Assumption (1.3), which provides a time dependent $\mathbf{L}^{\infty}$ bounds for the numerical solution. We choose a new approach, relying on the theory of discontinuous flux, see [AG2003, AJG2004, AGM2005, AKR2011, Diehl1995, GR1992, KRT2003, KT2004, Mishra2005, Towers2000] and the references therein. For the construction of the scheme, the space dependency of $H$ is discretized, and the idea is to treat each edge of the mesh as an interface of a discontinuous flux problem. Outside the compact of space dependency, our scheme reduces to a standard Godunov method. The contribution of the discontinuous flux theory is that we can build by hand discrete steady states of the scheme, making it, in a way well-balanced.
@@
2 Description of the scheme and main result
@@
$\forall j\in\mathbb{Z},\;u_{j}^{o}\coloneq\frac{1}{\Delta x}\int_{I_{j}}u_{o}(x)\mathinner{\mathrm{d}{x}},\quad\text{for a.e. }x\in\mathbb{R},\;\xi_{\Delta}(x)\coloneq\sum_{j\in\mathbb{Z}}x_{j}\mathbf{1}_{I_{j}}(x).$
Fix $n\in\mathbb{N}$. Let us explain how, given $(u_{j}^{n})_{j\in\mathbb{Z}}$, we determine $(u_{j}^{n+1})_{j\in\mathbb{Z}}$. Set $u^{n}\coloneq\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{I_{j}}$. Let us call $\mathcal{U}^{n}\in\mathbf{C}([t^{n},t^{n+1}],\mathbf{L}_{\mathbf{loc}}^{1}(\mathbb{R}))$ the unique entropy solution to following Cauchy problem, set on $\mathopen{]}t^{n},t^{n+1}\mathclose{[}\times\mathbb{R}$:
$\partial_{t}\mathcal{U}(t,x)+\partial_{x}\bigl(H(\xi_{\Delta}(x),\mathcal{U}(t,x))\bigr)$ $=0$
$\mathcal{U}(t^{n},x)$ $=u^{n}(x).$ (2.1)
We recall in Section 3.1 the meaning of solutions to conservation laws with discontinuous flux we use and also establish that (2) admits a unique solution, which justifies the definition of $\mathcal{U}^{n}$. Notice that since the space dependency is localized in a compact subset of $\mathbb{R}$, outside this compact, (2) reduces to a standard homogeneous conservation law. It then makes sense to $u_{j}^{n+1}$ as
$\forall j\in\mathbb{Z},\quad u_{j}^{n+1}\coloneq\frac{1}{\Delta x}\int_{I_{j}}\mathcal{U}^{n}(t^{n+1}-,x)\mathinner{\mathrm{d}{x}},$ (2.2)
where $\mathcal{U}^{n}(t^{n+1}-)\in\mathbf{L}_{\mathbf{loc}}^{1}(\mathbb{R})$ denotes the left-hand limit of $\mathcal{U}^{n}$ at $t=t^{n+1}$.
@@
Assume that $H$ satisfies (C3)–(CH)–(CVX) and fix $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. There exist two constants $\underline{u}$, $\overline{u}$ depending only on $H$ and $u_{o}$ such that under the CFL condition
$2\lambda L\leq 1,\quad\lambda\coloneq\frac{\Delta t}{\Delta x},\;L\coloneq\sup_{\underset{\underline{u}\leq p\leq\overline{u}}{x\in\mathbb{R}}}{\left|\partial_{u}H(x,p)\right|},$ (2.3)
@@
    $\text{for a.e.}\;(t,x)\in\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\quad\underline{u}\leq u_{\Delta}(t,x)\leq\overline{u}.$ (2.4)
@@
The proof of (M2) is then somewhat standard in the sense that if follows the classical steps of finite volume schemes analysis. All the details are presented in Section 3.3. The technical detail in the convergence proof is the use of the compensated compactness method [LuBook, SerreBook, Tartar1979, Tartar1982], since global $\mathbf{BV}$ bounds are not expected in the context of discontinuous flux. The genuine nonlinearity of $H$ under Assumption (CVX) is exploited.
Remark 2.1.
A byproduct of the analysis in that under the assumptions of Theorem 2.1, the sequence
$\rho_{\Delta}\coloneq\sum_{n=0}^{+\infty}\sum_{j\in\mathbb{Z}}u_{j}^{n}\mathbf{1}_{[t^{n},t^{n+1}\mathclose{[}\times I_{j}}$
also converges towards the entropy solution of (1), see Step 1 in Section 3.3. This sequence is meaningful since it is the one actually implemented in a computer.